show me step by step tou inherited $12,000 from my uncke Michael. If I invest it and average an 8% annual rate of return(compound annually), how much will I have in 20 years? which formula would I use? what is my compound factor?

FV formula? FVA?
Answer #1

Solution :- Future value = Present value * (1 + Rate of return)Time period in years

= 12000 * (1 + 0.08)20

= 12000 * (1.08)20

= 12000 * 4.661 (approx).

= $ 55932.

Conclusion :- In 20 years, You will have $ 55932 when you invest $ 12,000 at 20 years ago.